Pump A can unload a tanker in 15 hours, Pump B can unload the same tanker in 25 hours. Because of an approaching storm, both pumps are used to empty the tanker. If the pumps begin at 2:30 a.m., what time will the job be complete?
Answer by ptaylor(1894)

Let x=number of hours that it takes both pumps to unload all the tanker
Pump A unloads at the rate of 1/15th tanker per hour
Pump B unloads at the rate of 1/25th tanker per hour
Pump A + Pump B together unloads at the rate of 1/15+1/25 tanker per hour:
1/15=5/75
1/25=3/75
5/75+3/75=8/75 tanker per hour
So, Pump A + Pump B together unloads at the rate of 8/75 tanker per hour

Now we know that the rate that both unload times number of hours to unload the whole thing equals the whole thing off loaded. So our equation to solve is:

(8/75)x=75/75 (Note: 75/75 denotes all the tankers load)
multiply both sides by 75 and we get
8x=75
x=9.375 hours-9h 22.5m
And 2:30 am plus 9h 22.5m=11:52.5 AM------------------answer



Hope this helps----ptaylor